OpenStudy (anonymous):

solve differential equation: Q'(t)=-2Q(t)/435.475-t

OpenStudy (anonymous):

dQ/dt= -2Q ----------- 435.475-t 1/Q dq= 1/435.475-t dt

OpenStudy (anonymous):

integrate both sides

OpenStudy (anonymous):

please continue for me

OpenStudy (anonymous):

integral of 1/Q= ln Q +c

OpenStudy (anonymous):

my problem is solving the right side of the equation please continue